28:39 What is the x-coordinate of this solution?
2x+3y=-27
-3x+2y=23
My Work is Shown Below:
3(2x-3y=-27)
-3+2y=23
__________
6x-9y=81
-3x+2y=23
_________
-18x-7y=104
+18 +18
-7y=122
__ ___
-7 -7
y= -17.42
________________________________________________________________________________
Is that correct?

Answer by Alan3354(69443)

What is the x-coordinate of this solution?
2x+3y=-27
-3x+2y=23
My Work is Shown Below:
3(2x-3y=-27)
-3+2y=23
__________
6x-9y=81
-3x+2y=23
_________
-18x-7y=104
+18 +18
-7y=122 ********** What happened to the x? -18x + 18 <> 0
__ ___
-7 -7
y= -17.42
________________________________________________________________________________
Is that correct?
================
2x+3y=-27
-3x+2y=23
-----
6x + 9y = -81 eqn 1 times 3
-6x+ 4y = 46 eqn 2 times 2
------------- Add
13y = -35
y = -35/13